OpenStudy (anonymous):

Find the area of a circle whose area equals the half of the sum of the areas of two circles with radii 2 and 3. a) 13 b) 6.5 c) 5

OpenStudy (anonymous):

I'm pretty sure that it's B

OpenStudy (anonymous):

wow this is totally different thatn the last one

OpenStudy (anonymous):

It is/

OpenStudy (anonymous):

first lets compute the area of each circle the one with radius 2, and the one with radius 3

OpenStudy (anonymous):

or i guess i should as, "do you know those areas?"

OpenStudy (anonymous):

4 and 9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually \(4\pi\) and \(9\pi\) add them up and divide by 2

OpenStudy (anonymous):

so yeah, B although i assume there is a \(\pi\) there somewheres

OpenStudy (anonymous):

That's what I thought, but I guess i forgot to add pi

OpenStudy (anonymous):

i am sure it is in the answer choices right? \[6.5\pi\] not just \(6.5\)

OpenStudy (anonymous):

Yeah it doesn't show when I cut and pasted the questions. It did that with the last one too Lol!

OpenStudy (anonymous):

now i have a question who is your favorite trumpet player?

OpenStudy (anonymous):

Oh boy....

OpenStudy (anonymous):

Um in real life right now, it would have to be my previous band director

OpenStudy (anonymous):

Other than that I really can't decide between duke ellington and dizzy gelepsie

OpenStudy (anonymous):

oh ok i was going to pick Louis Armstrong Dizzy Gillespie Lee Morgan

OpenStudy (anonymous):

the duke did not pay trumpet though

OpenStudy (anonymous):

I thought he did for sure

OpenStudy (anonymous):

Oh well just a little misunderstanding I suppose

OpenStudy (anonymous):

check this when you get time Lee Morgan Sidewinder https://www.youtube.com/watch?v=lRIXN9f-Ap4

OpenStudy (anonymous):

But Dizzie Gelepsie was pretty Hillariouse to me. Especially how he puffed his cheeks out like a tuba when he played

OpenStudy (anonymous):

lol yeah, not to reveal my age, but i did see Dizzy once

OpenStudy (anonymous):

And yes I definitely will look it up

OpenStudy (anonymous):

Wait what! Really? That's so cool

OpenStudy (anonymous):

it was !

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

i was a youth , it was great

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I bet

OpenStudy (anonymous):

Wish I could've been

OpenStudy (anonymous):

got more math or are you done?

OpenStudy (anonymous):

Oh yeah I do

OpenStudy (anonymous):

ok lets see it !

OpenStudy (anonymous):

Just a couple more that's all

OpenStudy (anonymous):

hey, you should email me sometime

OpenStudy (anonymous):

satellite73.openstudy@gmail.com

OpenStudy (anonymous):

trumpetmagic2015@gmail.com

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Same

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Btw I'll send you some youtube links of the old band I used to play in

OpenStudy (anonymous):

great!